Microbiology difficult multiple choice questions I am unsure about!

1. eyrhtomycin acts on the protein synthesis of 70s ribosome high concentrations in human apllications would be detremental due to its influcence on which organelle of its eucaryotic cell
a. the ribosome
b. the nucleus
c. the lysosome
d. the golgi apparatus
e. the mitochondrian

2. if the average bacteria count is 296 and has a dilution factor of 1:10,000 the total number of bacteria per ml in the original sample would have been
a. 960,000

b.104
c.3,000,000
d.2,800
e.10,000

3.  if 0.1 ml of a culture is added to a dilution bottle containing 9.9ml of sterile water. the dilution factor for the contents would be
a.1:99
b.1:100
c.1:1000
d.1:10
e.1:1
Answer #1

eyrhtomycin acts on the protein synthesis of 70s ribosome high concentrations in human apllications would be detremental due to its influcence on which organelle of its eucaryotic cell the mitochondrian ( because mitochondria contain 70S subunit of ribosomes and it will affect by eyrhtomycin and disturb protein synthesis in mitochondria.)

if the average bacteria count is 296 and has a dilution factor of 1:10,000 the total number of bacteria per ml in the original sample would have been 3,000,000 ( total number of cells per ml can be count by number of cells in plate multiply by dilution factor = 296 * 10^4= 2960000 approximate equal to 3000000)

if 0.1 ml of a culture is added to a dilution bottle containing 9.9ml of sterile water. the dilution factor for the contents would be 1:100(dilution factor calculated by formula total volume/ aliquot volume = 0.1+9.9/0.1= 10/0.1=100= 1:100)